# AtomGraph: Tackling Atomicity Violation in Smart Contracts using Multimodal GCNs

### v1

@@ 1. Introduction
In recent years, frequent blockchain security incidents have posed severe challenges to the digital asset ecosystem. Critical infrastructures such as Decentralized Exchanges (DEX), Decentralized Finance (DeFi) protocols, and cross-chain bridges (gao2025implementation; shen2025blockchain) have repeatedly become targets of attacks, causing significant economic losses and trust crises. According to statistics from SlowMist Hacked (slowmist), the February 2025 theft on the Bybit platform was particularly noteworthy. Over \$1.46 billion was stolen, making it one of the largest cryptocurrency thefts within the past three years. Behind these massive losses lie the complex and severe security challenges that blockchain systems have long faced (chen2025chatgpt; gong2025information; zhang2025risk; xiang2025security). Among the top 10 common attack types, smart contract vulnerabilities dominate, accounting for as much as 26.55% (wu2025security; peng2025mining). Among these numerous security risks, atomicity violations represent a particularly insidious and destructive smart contract defect. They violate the fundamental principle that transactions must be fully successful or completely fail. Such defects are difficult to identify through manual audits and also pose significant challenges to traditional automated analysis tools.
To address these challenges, we propose AtomGraph, a novel automated detection framework for atomicity violations in smart contracts. This framework is based on multimodal GCNs and achieves deep analysis of bytecode through a collaborative learning paradigm for heterogeneous information (10646885). Specifically, AtomGraph first converts the contract into a CFG and learns features from two complementary modalities: (1) Structural Modality: capturing control flow dependencies from the topological structure of the CFG. (2) Semantic Modality: Understands the execution intent of the code from opcode sequences in basic blocks. To effectively fuse these heterogeneous features, we innovatively designed an adaptive weighted fusion mechanism that dynamically evaluates and adjusts modal weights, achieving optimal feature combination within a unified embedding space. Finally, the GCN is used to detect graph-level atomicity violations on contracts.
@@ 2. Related Work
Smart Contract Vulnerability Detection. These methods are primarily categorized into static and dynamic analysis. Static analysis can be further subdivided into symbolic execution, such as Oyente (Luu_2016_Making) and Slither (Feist_2019_Slither). They also include pattern-matching tools like Securify (tsankov2018securify) and Ponzi scheme analysis (zhang2025security). Other methods such as taint analysis with Ethainter (brent2020ethainter) or intermediate representation analysis using Vandal (brent2018vandal), Conkas (Veloso__Conkas) and MoveScanner (luo2025movescanner). These methods typically offer advantages in execution speed and high coverage, but they also commonly suffer from high False Positive Rates (FPRs). Dynamic analysis methods, including Echidna (grieco2020echidna) and Harvey (wustholz2020harvey), detect vulnerabilities by executing code and capturing runtime issues (li2025penetrating; li2025interaction). However, they face limitations due to path explosion and difficulty generating test cases. Additionally, hybrid approaches like Mythril (mythril) and Manticore (Mossberg_2019_Manticore) combine static and dynamic analysis. They enhance detection accuracy through mutual feedback, but they generally require more computational resources and time.
Deep learning-based methods demonstrate significant potential (Zhuang_2020_Smart; ZHU2021107920; peng2025multicfv; wang2025ai; li2025facial). Tools such as Clear (chen2024improving) and SCVHunter (luo2024scvhunter) leverage models to learn vulnerability patterns and Features from large-scale smart contract samples, enabling automated detection and prediction of novel vulnerabilities.
Smart Contract Atomicity Violations. The atomicity principle requires that a transaction must possess an all-or-nothing (zhang2025bounded). Either all components execute successfully, or all must be rolled back and canceled, preventing any intermediate or partially completed states (li2024detecting). In smart contracts, when a sequence of operations that must be executed as a whole is unexpectedly interrupted or only partially completed, it results in an atomicity violation, leading to inconsistent states (liu2025detecting).
Common forms of atomicity violations in smart contracts include: (1) Reentrancy attacks: Malicious re-entry of external calls before state updates complete (song2025silence). (2) Check-Effect-Interaction race conditions: Exploitation of the time window between check conditions and resource usage. (3) Improper Exception Handling: Partial state modifications occur during exceptions without proper rollback. (4) Cross-Domain Non-Atomicity: Partial failures across multiple smart contracts or different blockchain systems (liu2025empirical; huang2025comparative).
@@ 3. Method
CFG Construction. The CFG is a structured representation of smart contract execution logic, clearly illustrating the basic blocks, branching structures, and control flow transfers within a contract (zhen2024gnn). Building upon existing research, we employ static analysis methods to achieve precise conversion from opcodes to CFGs. First, we disassembled the raw bytecode stream into an opcode sequence that complies with the Ethereum Virtual Machine (EVM) (ma2025uncovering)specification.
Based on disassembled opcodes, we utilize standard basic block partitioning algorithms to identify code structures. Specifically, a basic block is a contiguous sequence of instructions with a single entry point and a unique exit point (wang2023smart). Basic block boundaries are defined by specific control flow instructions such as JUMP, JUMPI, STOP, REVERT, RETURN, and the target addresses of jump instructions. After identifying all basic blocks, we construct the edges of the CFG by analyzing jump relationships between blocks (10.1145/3699597). For example, when a basic block ends with a conditional jump instruction (JUMPI), two outgoing edges are generated: one pointing to the basic block at the jump target address, and another pointing to the next basic block in sequential execution. By systematically connecting all basic blocks, we construct a complete CFG that precisely reflects the underlying execution logic of the contract.
@@
Graph Convolutional Network. After obtaining high-quality multimodal node features, we employ GCN as the core classification model, integrating node neighborhood information through multiple GraphConv layers (shang2025cegt). The fundamental propagation rule for each layer is as follows:
@@
The AdamW optimizer is employed with a dynamic learning rate scheduling strategy (10480574). We integrate focus loss functions, dynamic class weighting, and data augmentation techniques to address class imbalance. Dropout regularization and gradient clipping enhance model generalization and training stability.
@@ 4. Experiments
Dataset. This study’s dataset is based on the SmartBugs Wild (DurieuxEtAl2020ICSE) collection of 47,398 real contracts and DAppSCAN (10486822), which includes 21,457 contracts and 608 audit reports. A specialized dataset comprising 1,880 smart contract samples was constructed through systematic screening and manual annotation. We used 90% of the dataset for training and 10% for testing. Table 1 summarizes the smart contract data used.
@@
Answer to RQ2: Comparative Analysis. As shown in Table 3, we conducted comparative experiments of AtomGraph with other 10 smart contract analysis tools. Symbolic execution tools Mythril, Osiris (torres2018osiris), Oyente, and Pakala (pakala) exhibit the typical characteristics of ”high precision, low recall”. Oyente achieves 100% precision but only 4.40% recall, while Osiris achieves 96.67% precision but only 18.45% recall. This reflects the path explosion problem of symbolic execution when dealing with complex state spaces, leading to insufficient analysis coverage. Pattern-matching tools Securify are similarly constrained by the scope of predefined rules, achieving 97.20% precision but only 14.32% recall. Taint analysis tools Ethainter and intermediate representation analysis tools Vandal show the opposite trend: Vandal relaxes detection conditions to improve recall 44.95%, but this introduces a high FPR of 40%.
Additionally, tools like Conkas and Ethor (schneidewind2020ethor) achieved moderate F1 scores, with Conkas scoring 43.73% and Ethor scoring 40.38%. While HoneyBadger (torres2019art), which focuses on honeypot detection through symbolic execution expansion, achieves high precision but is limited by its narrow detection scope, resulting in an overall recall rate of only 24.04%. AtomGraph’s multimodal architecture overcomes these limitations by integrating structural and semantic modalities, achieving an F1 score of 96.97% while maintaining a FPR of 0.05%.

### v2

@@ 1. Introduction
In recent years, frequent blockchain security incidents have posed severe challenges to the digital asset ecosystem. Critical infrastructures such as Decentralized Exchanges (DEX), Decentralized Finance (DeFi) protocols, and cross-chain bridges (Gao et al., 2025; Shen et al., 2025; Huang et al., 2024; Aguilar et al., 2024) have repeatedly become targets of attacks, causing significant economic losses and trust crises. According to statistics from SlowMist Hacked (Zone, 2025), the February 2025 theft on the Bybit platform was particularly noteworthy. Over \$1.46 billion was stolen, making it one of the largest cryptocurrency thefts within the past three years. Behind these massive losses lie the complex and severe security challenges that blockchain systems have long faced (Chen et al., 2025; Gong et al., 2025; Zhang et al., 2025b; Xiang et al., 2025b). Among the top 10 common attack types, smart contract vulnerabilities dominate, accounting for as much as 26.55% (Wu et al., 2025b; Peng et al., 2025a). Among these numerous security risks, atomicity violations represent a particularly insidious and destructive smart contract defect. They violate the fundamental principle that transactions must be fully successful or completely fail. Such defects are difficult to identify through manual audits and also pose significant challenges to traditional automated analysis tools.
To address these challenges, we propose AtomGraph, a novel automated detection framework for atomicity violations in smart contracts. This framework is based on multimodal GCNs and achieves deep analysis of bytecode through a collaborative learning paradigm for heterogeneous information (Wang et al., 2024a). Specifically, AtomGraph first converts the contract into a CFG and learns features from two complementary modalities: (1) Structural Modality: capturing control flow dependencies from the topological structure of the CFG. (2) Semantic Modality: Understands the execution intent of the code from opcode sequences in basic blocks. To effectively fuse these heterogeneous features, we innovatively designed an adaptive weighted fusion mechanism that dynamically evaluates and adjusts modal weights, achieving optimal feature combination within a unified embedding space. Finally, the GCN is used to detect graph-level atomicity violations on contracts.
@@ 2. Related Work
Smart Contract Vulnerability Detection. These methods are primarily categorized into static and dynamic analysis. Static analysis can be further subdivided into symbolic execution, such as Oyente (Luu et al., 2016) and Slither (Feist et al., 2019). They also include pattern-matching tools like Securify (Tsankov et al., 2018) and Ponzi scheme analysis (Zhang et al., 2025a). Other methods such as taint analysis with Ethainter (Brent et al., 2020) or intermediate representation analysis using Vandal (Brent et al., 2018), Conkas (Veloso, 2023) and MoveScanner (Luo et al., 2025; Wang and Zhao, 2024; Wang et al., 2024b). These methods typically offer advantages in execution speed and high coverage, but they also commonly suffer from high False Positive Rates (FPRs). Dynamic analysis methods, including Echidna (Grieco et al., 2020) and Harvey (Wüstholz and Christakis, 2020), detect vulnerabilities by executing code and capturing runtime issues (Li et al., 2025b, a, 2024b). However, they face limitations due to path explosion and difficulty generating test cases. Additionally, hybrid approaches like Mythril (Mueller, 2025) and Manticore (Mossberg et al., 2019) combine static and dynamic analysis. They enhance detection accuracy through mutual feedback, but they generally require more computational resources and time.
Deep learning-based methods demonstrate significant potential (Zhuang et al., 2021; Zhu et al., 2021; Peng et al., 2025b; Wang and Li, 2025; Li et al., 2025c). Tools such as Clear (Chen et al., 2024a) and SCVHunter (Luo et al., 2024) leverage models to learn vulnerability patterns and Features from large-scale smart contract samples, enabling automated detection and prediction of novel vulnerabilities.
Smart Contract Atomicity Violations. The atomicity principle requires that a transaction must possess an all-or-nothing (Zhang et al., 2025c). Either all components execute successfully, or all must be rolled back and canceled, preventing any intermediate or partially completed states (Li et al., 2024a). In smart contracts, when a sequence of operations that must be executed as a whole is unexpectedly interrupted or only partially completed, it results in an atomicity violation, leading to inconsistent states (Liu et al., 2025b; Wu et al., 2025a; Xiao et al., 2025; Cai et al., 2024).
Common forms of atomicity violations in smart contracts include: (1) Reentrancy attacks: Malicious re-entry of external calls before state updates complete (Song et al., 2025). (2) Check-Effect-Interaction race conditions: Exploitation of the time window between check conditions and resource usage. (3) Improper Exception Handling: Partial state modifications occur during exceptions without proper rollback. (4) Cross-Domain Non-Atomicity: Partial failures across multiple smart contracts or different blockchain systems (Liu et al., 2025a; Huang et al., 2025; Han et al., 2026; Chen et al., 2024b).
@@ 3. Method
CFG Construction. The CFG is a structured representation of smart contract execution logic, clearly illustrating the basic blocks, branching structures, and control flow transfers within a contract (Zhen et al., 2024; Arceri et al., 2024). Building upon existing research, we employ static analysis methods to achieve precise conversion from opcodes to CFGs. First, we disassembled the raw bytecode stream into an opcode sequence that complies with the Ethereum Virtual Machine (EVM) (Ma et al., 2025)specification.
Based on disassembled opcodes, we utilize standard basic block partitioning algorithms to identify code structures. Specifically, a basic block is a contiguous sequence of instructions with a single entry point and a unique exit point (Wang et al., 2023). Basic block boundaries are defined by specific control flow instructions such as JUMP, JUMPI, STOP, REVERT, RETURN, and the target addresses of jump instructions. After identifying all basic blocks, we construct the edges of the CFG by analyzing jump relationships between blocks (Xiang et al., 2025a). For example, when a basic block ends with a conditional jump instruction (JUMPI), two outgoing edges are generated: one pointing to the basic block at the jump target address, and another pointing to the next basic block in sequential execution. By systematically connecting all basic blocks, we construct a complete CFG that precisely reflects the underlying execution logic of the contract.
@@
Graph Convolutional Network. After obtaining high-quality multimodal node features, we employ GCN as the core classification model, integrating node neighborhood information through multiple GraphConv layers (Shang et al., 2025). The fundamental propagation rule for each layer is as follows:
@@
The AdamW optimizer is employed with a dynamic learning rate scheduling strategy (Zhou et al., 2024). We integrate focus loss functions, dynamic class weighting, and data augmentation techniques to address class imbalance. Dropout regularization and gradient clipping enhance model generalization and training stability.
@@ 4. Experiments
Dataset. This study’s dataset is based on the SmartBugs Wild (Durieux et al., 2020) collection of 47,398 real contracts and DAppSCAN (Zheng et al., 2024; Ayub et al., 2024), which includes 21,457 contracts and 608 audit reports. A specialized dataset comprising 1,880 smart contract samples was constructed through systematic screening and manual annotation. We used 90% of the dataset for training and 10% for testing. Table 1 summarizes the smart contract data used.
@@
Answer to RQ2: Comparative Analysis. As shown in Table 3, we conducted comparative experiments of AtomGraph with other 10 smart contract analysis tools (He et al., 2025; Sun et al., 2025). Symbolic execution tools Mythril, Osiris (Torres et al., 2018), Oyente, and Pakala (Palkeo, 2025) exhibit the typical characteristics of ”high precision, low recall”. Oyente achieves 100% precision but only 4.40% recall, while Osiris achieves 96.67% precision but only 18.45% recall. This reflects the path explosion problem of symbolic execution when dealing with complex state spaces, leading to insufficient analysis coverage. Pattern-matching tools Securify are similarly constrained by the scope of predefined rules, achieving 97.20% precision but only 14.32% recall. Taint analysis tools Ethainter and intermediate representation analysis tools Vandal show the opposite trend: Vandal relaxes detection conditions to improve recall 44.95%, but this introduces a high FPR of 40%.
Additionally, tools like Conkas and Ethor (Schneidewind et al., 2020) achieved moderate F1 scores, with Conkas scoring 43.73% and Ethor scoring 40.38%. While HoneyBadger (Torres et al., 2019), which focuses on honeypot detection through symbolic execution expansion, achieves high precision but is limited by its narrow detection scope, resulting in an overall recall rate of only 24.04%. AtomGraph’s multimodal architecture overcomes these limitations by integrating structural and semantic modalities, achieving an F1 score of 96.97% while maintaining a FPR of 0.05%.